OpenStudy (anonymous):

what is the inverse of h(x) = x3 + 5

OpenStudy (anonymous):

y=x^3+5

OpenStudy (anonymous):

now switch the x and y

OpenStudy (anonymous):

x=y^3+5

OpenStudy (anonymous):

now solve for y

OpenStudy (anonymous):

Can you run me through how to?

OpenStudy (anonymous):

I'm not sure how to do it

OpenStudy (anonymous):

ok, first subtract 5 from both sides

OpenStudy (anonymous):

You just get y^3

OpenStudy (anonymous):

i'm really bad at this

OpenStudy (anonymous):

nope, x-5 = y^3

OpenStudy (anonymous):

now take the cube root of both sides

OpenStudy (anonymous):

(x-5)^(1/3) = y

OpenStudy (anonymous):

That's the answer

OpenStudy (anonymous):

Thank you!